OpenStudy (anonymous):

Solve each equation. 8x^2-24=3x^2-9

OpenStudy (anonymous):

As well as Solve each equation. 0.2x2 + 1.2 = 3.4

OpenStudy (acxbox22):

subtract 3x^2 from both sides 5x^2=15 -------- 5 x^2=3 take the square root of both sides and round

OpenStudy (yanasidlinskiy):

Simplifying 8x^2 + -24 = 3x^2 + -9 Reorder the terms: -24 + 8x^2 = 3x^2 + -9 Reorder the terms: -24 + 8x^2 = -9 + 3x^2 Solving: -24 + 8x^2 = -9 + 3x^2 Solving for variable 'x'. Move all terms containing x to the left, all other terms to the right. Add '-3x^2' to each side of the equation. -24 + 8x^2 + -3x^2 = -9 + 3x^2 + -3x^2 Combine like terms: 8x^2 + -3x^2 = 5x^2 -24 + 5x^2 = -9 + 3x^2 + -3x^2 Combine like terms: 3x^2 + -3x^2 = 0 -24 + 5x^2 = -9 + 0 -24 + 5x^2 = -9 Add '24' to each side of the equation. -24 + 24 + 5x2 = -9 + 24 Combine like terms: -24 + 24 = 0 0 + 5x^2 = -9 + 24 5x^2 = -9 + 24 Combine like terms: -9 + 24 = 15 5x^2 = 15 Divide each side by '5'. x^2 = 3 Simplifying x^2 = 3 Take the square root of each side: x = {-1.732050808, 1.732050808}